# Robustness is Important: Limitations of LLMs for Data Fitting

### v2

@@ Abstract
Large Language Models (LLMs) are being applied in a wide array of settings, well beyond the typical language-oriented use cases. In particular, LLMs are increasingly used as a plug-and-play method for fitting data and generating predictions. Prior work has shown that LLMs, via in-context learning or supervised fine-tuning, can perform competitively with many tabular supervised learning techniques in terms of predictive performance. However, we identify a critical vulnerability of using LLMs for data fitting – making changes to data representation that are completely irrelevant to the underlying learning task can drastically alter LLMs’ predictions on the same data. For example, simply changing variable names can sway the size of prediction error by as much as 82% in certain settings. Such prediction sensitivity with respect to task-irrelevant variations manifests under both in-context learning and supervised fine-tuning, for both close-weight and open-weight general-purpose LLMs. Moreover, by examining the attention scores of an open-weight LLM, we discover a non-uniform attention pattern: training examples and variable names/values which happen to occupy certain positions in the prompt receive more attention when output tokens are generated, even though different positions are expected to receive roughly the same attention. This partially explains the sensitivity in the presence of task-irrelevant variations. We also consider a state-of-the-art tabular foundation model (TabPFN) trained specifically for data fitting. Despite being explicitly designed to achieve prediction robustness, TabPFN is still not immune to task-irrelevant variations. Overall, despite LLMs’ impressive predictive capabilities, currently they lack even the basic level of robustness to be used as a principled data-fitting tool.
@@ 1 Introduction
Large language models (LLMs), as one of the most representative Generative AI applications, have been adopted across a wide variety of domains and tasks. While the majority of LLMs’ use cases are expectedly language-related (translation, Q&A, writing, ideation, etc.), there are also increasing uses of LLMs for tasks not apparently associated with language. We focus on one such practice, where LLM is used as an “off-the-shelf” data-fitting tool to generate predictions of interest. Data fitting refers to the task of using sets $\{(\boldsymbol{X},Y)\}$ of numerical input values $\boldsymbol{X}$ and target values $Y$ to learn (fit) a model $f(.)$ – i.e., $Y\sim f(\boldsymbol{X})$ – that can generate accurate predictions given inputs. For example, LLMs have been used for predicting property prices based on property features (e.g., number of bedrooms, bathrooms, etc.) and location information Tanlamai et al., (2024); performing time series forecasting and numeric prediction tasks in general by treating input data as a sequence of string tokens Gruver et al., (2023); Jin et al., (2023); Das et al., (2024); Goswami et al., (2024); Vacareanu et al., (2024); detecting anomalies in time series data Alnegheimish et al., (2024); Dong et al., (2024); Liu et al., (2025); generating synthetic data in data-hungry domains Seedat et al., (2023); Wang et al., (2024); Brand et al., (2023); Isomura et al., (2024); and support causal effect estimation by predicting counterfactual outcomes (Huynh et al.,, 2025). In all of these examples, LLMs essentially act as a drop-in replacement for traditional, tabular supervised learning techniques to generate predictions based on input data (See Appendix A for more details about these examples).
@@
A priori, it is unclear whether LLMs should be used for data fitting and predictions. On one hand, hailed as a general-purpose technology Eloundou et al., (2024), there is evidence that LLMs’ impressive capabilities extend beyond language tasks and can perform data fitting quite well. Prior work has shown that LLMs, with just a few input-output examples included in the prompt (i.e., using few-shot prompting), can outperform many tabular supervised learning techniques for regression tasks (Vacareanu et al.,, 2024). Furthermore, when feature names are semantically meaningful, LLMs can perform feature selection automatically, which may further benefit data fitting (Jeong et al.,, 2024). On the other hand, LLMs vs. tabular supervised learning techniques perform “data fitting” and “prediction” in fundamentally different ways. Tabular supervised learning techniques perform data fitting by following well-defined algorithmic procedures that fit parametric functions (e.g., linear regression) or non-parametric functions (e.g., random forest) to training data. The trained models then produce predictions by directly applying the learned function (i.e., plug in the feature values of new input data). In contrast, the vast majority of LLMs today are based on the transformer architecture, which auto-regressively generates the next token based on prior tokens. With this architecture, LLMs perform data fitting by (i) having training examples in the prompt, namely in-context learning Brown et al., (2020), or by (ii) supervised fine-tuning using the training data Hu et al., (2021). [^1] We do not consider the use case where LLMs generate and/or execute code to build tabular models. In such case, LLMs are being used as a coding tool rather than a data-fitting tool. They generate predictions one token at a time, where each token may correspond to only a segment of the prediction value. As an illustrative example, consider how an LLM generates a numerical prediction value. To generate a prediction value of 0.1875, OpenAI’s GPT-4o model needs to generate 4 tokens precisely in the following order: “0”, “.”, “187”, and “5”. [^2] This example can be replicated at: https://platform.openai.com/tokenizer These fundamental differences raise the question whether LLMs provide a principled or robust way to fit numerical data and obtain predictions.
@@
However, just because LLMs can be used for data fitting does not mean they should be. A principled data-fitting technique not only needs to achieve reasonable predictive performance, but also needs to possess at least a basic level of robustness so as not to be affected by factors that are fundamentally unrelated to the task itself. Our analyses show that LLMs have alarmingly poor prediction robustness. In particular, we apply several different types of task-irrelevant variations – such as altering variable names, variable (column) order, example (row) order, and data format – none of which has any effect on the underlying data-fitting task. While canonical tabular supervised learning techniques are completely unaffected (essentially by their design), LLMs’ predictions for the same data can change dramatically, causing the size of prediction error to sway as much as 82% in some settings. Further statistical tests confirm that such prediction sensitivity is indeed attributable to the task-irrelevant variations and cannot be explained by randomness in the LLM token generation process.
@@
Our results reveal a fundamental challenge of using LLMs for data fitting: prediction robustness with respect to task-irrelevant variations, which is completely taken for granted when using tabular supervised learning techniques, can hardly be guaranteed in LLMs; at best, it can only be partially engineered in special-purpose foundation models. As an analogy, using LLMs for data fitting currently can be viewed like using a calculator for adding numbers, where the calculator produces a different sum when given the same set of numbers in a different order. Therefore, we caution against the increasingly prevalent use of LLMs as a black-box data-fitting tool – the prediction quality cannot be ensured due to lack of robustness with respect to task-irrelevant variations.
Beyond data fitting, it is worth highlighting that our findings also have broader implications for current LLMs as a paradigm of learning. We take the data-fitting task not as the ultimate objective per se but as an opportunity to better understand the limitations of LLMs. The ability to disentangle what is relevant vs. irrelevant for a task is widely recognized as a fundamental requirement for forming abstraction (Giere,, 2010; Weisberg,, 2012), learning conceptual categories (Rosch,, 2024), cognitive processing (e.g., the selective attention theory Broadbent,, 1958; Johnston and Dark,, 1986), all of which are important in problem-solving. When LLMs change their predictions in response to task-irrelevant variations, it calls into question whether LLMs possess basic “competence” for the data-fitting task. Together with recent work showing that inconsequential changes in prompting can significantly affect LLMs’ performance on solving math problems (Mirzadeh et al.,, 2024) and imitating human behaviors (Gao et al.,, 2025), we highlight that the use of current LLMs as principled problem-solving tools requires careful reconsideration.
@@ 2 Materials and Methods
Because modern LLMs have been pre-trained with massive datasets collected from the Internet, it is inappropriate to evaluate LLMs’ data-fitting capabilities on publicly available datasets which may have already been exposed to the LLMs during training. Therefore, we create synthetic data from scratch for all experiments in this paper. Specifically, we simulate the following linear data generation process for $p$ input variables:
@@
In the baseline setting, $p=10$ and each input variable $X_{i}$ is randomly and independently sampled from either $N(0,1)$ or $\exp(1)$; coefficients $\{\beta_{1},\ldots,\beta_{p}\}$ are sampled from $Unif[0,1]$ and fixed; and $\varepsilon$ is randomly sampled from $N(0,0.1^{2})$. We choose such a simple linear data-generation process for two reasons. First, its parametric form gives us fine-grained control over different aspects of the data-fitting task, such as number of input variables ($p$) and distributional characteristics of these variables. Second, it represents an “easy” data-fitting task. If LLMs exhibit problematic data-fitting capabilities under this setup, it is more likely due to a fundamental limitation of current LLMs rather than difficulty of the data-fitting tasks. We use this data-generation process to simulate a training dataset of 4,000 instances and a testing dataset of 1,000 instances.
@@
We focus on two specific data-fitting capabilities. We first evaluate the performance level, i.e., how well LLMs fit data and generate predictions compared to tabular supervised learning techniques. This determines whether LLMs have the capability to fit data accurately. More importantly, we also evaluate prediction sensitivity, i.e., whether LLM-generated predictions are robust with respect to changes in the inputs that are irrelevant to the data-fitting tasks (and are extraneous for tabular supervised learning techniques). This sheds light on whether LLMs should be used for data fitting.
@@
We carry out comprehensive data-fitting experiments on GPT-4o-mini (a representative general-purpose close-weight LLM), [^3] Among many close-weight commercial LLMs, GPT-4o-mini is one of the most cost-effective flagship LLMs that support both completion and fine-tuning. Llama-3-8B-instruct (a representative general-purpose open-weight LLM), as well as TabPFN (currently best-performing tabular foundation model specifically trained for data-fitting tasks, Hollmann et al.,, 2025). We perform LLM-based data fitting via both in-context learning (ICL) and supervised fine-tuning (SFT). For GPT-4o-mini, we use OpenAI’s batch processing API for ICL and fine-tuning API for SFT. For Llama-3-8B-instruct, we download the model from Hugging Face [^4] Source: https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct. and carry out ICL using resources provided by Minnesota Supercomputing Institute. For the tabular foundation model (TabPFN), we follow TabPFN documentation to perform ICL using its Python implementation. The detailed prompts and procedures for both ICL and SFT are reported in Appendix B. Given the objective nature of the task, we use greedy decoding for all LLMs (i.e., selecting the next token as the one receiving highest predicted probability). This is done by setting LLM generation temperature to 0 and disabling random token sampling whenever possible. We also fix the random seed during SFT of LLMs as well as the training of all tabular supervised learning models to ensure reproducibility and fair comparisons across different models. All models are evaluated on the same 1,000 testing instances. Replication materials can be accessed at https://github.com/heziiiiiiiii/LLMs-for-data-fitting.
@@ 3.1 Performance Level
Starting with GPT-4o-mini, we evaluate the LLMs’ predictive performance level as compared to six widely-used tabular supervised learning techniques, including linear regression (the oracle model in this case), LASSO regression, $k$-NN, Support Vector Regression (SVR), Random Forest, and Multi-Layer Perception (MLP). Each supervised technique is trained on the same training data of 4,000 instances, with hyper-parameters tuned via 5-fold cross-validation. For reference, we also include a naive benchmark, namely using the mean of outcome variable on the training data as (constant) prediction. For GPT-4o-mini under SFT, we use all 4,000 training instances for fine-tuning. Under ICL, following the common practice in few-shot prompting, we include either 10 or 20 instances sampled from training data into the task prompt (denoted as ICL10 and ICL20). We also explore an ICL prompt with 500 instances (denoted as ICL500), representing a large-scale setting that is still within the context window of GPT-4o-mini. We evaluate predictive performance on the testing data using the standard Mean Absolute Error (MAE) metric. The results are shown in Table 1.
@@
Not surprisingly, linear regression and other tabular supervised learning techniques that can sufficiently approximate a linear relationship (LASSO, SVR, and MLP) all perform quite well on this task, whereas Random Forest and $k$-NN perform comparatively worse. More importantly, LLM-based data fitting achieves respectable performance. With SFT, LLM outperforms Random Forest and $k$-NN and catches up with the performance of the other techniques. With ICL, even with only 10 or 20 examples (substantially smaller “training” size than the other models), it still meaningfully outperforms the naive benchmark, indicating that LLM can conduct few-shot learning to some degree in a prediction task. Further increasing the number of examples in ICL to 500 does not necessarily improve predictive performance. [^5] We further consider a few variations to the above baseline setting, including different numbers of variables, variable scales, and the presence of outliers in data. Overall, we find that LLMs achieve reasonable performance levels across these variations, especially with SFT.

### v3

@@ Abstract
Large Language Models (LLMs) are being applied in a wide array of settings, well beyond typical language-oriented use cases. In particular, LLMs are increasingly used as a plug-and-play method for fitting data and generating predictions. Prior work has shown that LLMs, via in-context learning or supervised fine-tuning, can perform competitively with many tabular supervised learning techniques on predictive performance. However, we identify a critical vulnerability of using LLMs for data fitting – making changes to data representation that are completely irrelevant to the underlying learning task can drastically alter LLMs’ predictions on the same data. For example, simply changing variable names can sway the size of prediction error by as much as 82% in certain settings. Such prediction sensitivity with respect to task-irrelevant variations manifests under both in-context learning and supervised fine-tuning, for both close-weight and open-weight general-purpose LLMs. Moreover, by examining the attention scores of an open-weight LLM, we discover a non-uniform attention pattern: training examples and variable names/values occupying certain positions in the prompt receive more attention when generating output tokens, even though fundamentally there should not be different emphasis a priori on specific data rows / columns. This partially explains the sensitivity due to task-irrelevant variations. We also consider a state-of-the-art tabular foundation model (TabPFN) trained specifically for data fitting. Despite being explicitly designed to achieve prediction robustness, TabPFN is still not immune to task-irrelevant variations. Overall, despite LLMs’ impressive predictive capabilities, currently they lack even the basic level of robustness to be used as a principled data-fitting tool.
@@ 1 Introduction
Large language models (LLMs), as one of the most representative Generative AI applications, have been adopted across a wide variety of domains and tasks. While the majority of LLMs’ use cases are expectedly language-related (translation, Q&A, writing, ideation, etc.), there are also increasing uses of LLMs for tasks not apparently associated with language. We focus on one such practice, where LLM is used as an “off-the-shelf” data-fitting tool to generate predictions of interest. Data fitting refers to the task of using sets $\{(\boldsymbol{X},Y)\}$ of numerical input values $\boldsymbol{X}$ and target values $Y$ to learn (fit) a model $f(.)$ – i.e., $Y\sim f(\boldsymbol{X})$ – that can generate accurate predictions given inputs. For example, LLMs have been used for predicting property prices based on property features (e.g., number of bedrooms, bathrooms, etc.) and location information Tanlamai et al., (2024); performing time series forecasting and numeric prediction tasks in general by treating input data as a sequence of string tokens Gruver et al., (2023); Jin et al., (2023); Das et al., (2024); Goswami et al., (2024); Vacareanu et al., (2024); detecting anomalies in time series data Alnegheimish et al., (2024); Dong et al., (2024); Liu et al., (2025); generating synthetic data in data-hungry domains Seedat et al., (2023); Wang et al., (2024); Brand et al., (2023); Isomura et al., (2024); supporting causal effect estimation by predicting counterfactual outcomes (Huynh et al.,, 2025); and predicting operational metrics in complex systems (Akhauri et al.,, 2025). Essentially, LLMs act as a drop-in replacement for traditional, tabular supervised learning techniques to generate predictions based on input data (See SI Appendix Table S1 for more details about these examples). Importantly, in all of these examples, LLMs are not being used as a coding agent to simply generate and/or execute code that applies tabular supervised learning techniques; instead, LLMs act as the data-fitting tool itself.
@@
A priori, it is unclear whether LLMs should be used for data fitting and predictions. On one hand, hailed as a general-purpose technology Eloundou et al., (2024), there is evidence that LLMs’ impressive capabilities extend beyond language tasks and can perform data fitting quite well. Prior work has shown that LLMs, with just a few input-output examples included in the prompt (i.e., using few-shot prompting), can outperform many tabular supervised learning techniques for regression tasks (Vacareanu et al.,, 2024). Furthermore, LLMs can select features based on semantic meaning of feature names, which may further benefit data fitting (Jeong et al.,, 2024). On the other hand, LLMs vs. tabular supervised learning techniques perform “data fitting” and “prediction” in fundamentally different ways. Tabular supervised learning techniques perform data fitting by following well-defined algorithmic procedures that fit parametric functions (e.g., linear regression) or non-parametric functions (e.g., random forest) to training data. The trained models then produce predictions by directly applying the learned function (i.e., plug in the feature values of new input data). In contrast, the vast majority of LLMs today are based on the transformer architecture, which auto-regressively generates the next token based on prior tokens. With this architecture, LLMs perform data fitting by (i) having training examples in the prompt, namely in-context learning Brown et al., (2020), or by (ii) supervised fine-tuning using the training data Hu et al., (2021). [^1] We do not consider the use case where LLMs generate and/or execute code to build predictive models because, as discussed before, we observe an increasingly prevalent use of LLMs acting as a data-fitting tool itself, and that is the focus of our study. They generate predictions one token at a time, where each token may correspond to only a textual segment of the prediction value. As an illustrative example, consider how an LLM generates a numerical prediction value. To generate a prediction value of 0.1875, OpenAI’s GPT-4o model needs to generate 4 tokens precisely in the following order: “0”, “.”, “187”, and “5”. [^2] This example can be replicated at: https://platform.openai.com/tokenizer These fundamental differences raise the question whether LLMs provide a principled or robust way to fit numerical data and obtain predictions.
@@
However, just because LLMs can be used for data fitting does not mean they should be. A principled data-fitting technique not only needs to achieve reasonable predictive performance, but also needs to possess at least a basic level of robustness so as not to be affected by factors that are fundamentally unrelated to the task itself. Our analyses show that LLMs have alarmingly poor prediction robustness. In particular, we apply several different types of task-irrelevant variations – such as altering variable names, variable (column) order, example (row) order, and data format – none of which technically should have any effect on the underlying data-fitting task. Canonical tabular supervised learning techniques are, essentially by their design, either completely unaffected or affected to a minor degree explained by randomness (e.g., due to different random seeds that control certain stochastic aspects of the learning procedure). In contrast, LLMs’ predictions for the same data can change dramatically, causing the size of prediction error to sway as much as 82% in some settings. Further statistical tests confirm that such prediction sensitivity is indeed attributable to the task-irrelevant variations and cannot be explained by randomness in the LLM token generation process.
@@
Our results reveal a fundamental challenge of using LLMs for data fitting: prediction robustness with respect to task-irrelevant variations, which is generally taken for granted when using tabular supervised learning techniques, can hardly be guaranteed in LLMs; at best, it can only be partially engineered in special-purpose foundation models. As an analogy, using LLMs for data fitting currently can be viewed like using a calculator for adding numbers, where the calculator produces a substantially different sum when given the same exact set of numbers but in a different order. Thus, we caution against the increasingly prevalent use of LLMs as a black-box data-fitting tool – prediction quality cannot be ensured due to lack of robustness with respect to task-irrelevant variations.
It is crucial to note that the focus of our study is not on improving LLMs’ data-fitting and prediction capabilities per se. If one takes a very narrow focus on predictive performance and blissfully ignores anything else, one could potentially lean on the increasingly powerful tabular foundation models, or even an “ensemble” of LLMs (e.g., generate multiple predictions with varying LLMs/prompts and then take an average). However, we maintain the position that data-fitting is fundamentally about “learning from data”, from which good predictive performance emerges as a result (Vapnik,, 2013). As such, we use the data-fitting task as an important lens through which we seek to better understand the limitations of current LLMs as a paradigm of learning. The ability to disentangle what is relevant vs. irrelevant for a task is widely recognized as a fundamental requirement for forming abstraction (Giere,, 2010; Weisberg,, 2012), learning conceptual categories (Rosch,, 2024), cognitive processing (e.g., the selective attention theory Broadbent,, 1958; Johnston and Dark,, 1986), all of which are important in problem solving. When LLMs change their predictions in response to task-irrelevant variations, it raises the question whether LLMs possess basic “competence” for learning from data. Together with recent work showing that inconsequential changes in prompting can significantly affect LLM performance in solving math problems (Mirzadeh et al.,, 2024) and imitating human behaviors (Gao et al.,, 2025), we highlight that the use of current LLMs as principled learning and problem-solving tools requires careful reconsideration.
@@ 2 Materials and Methods
Because modern LLMs have been trained with massive datasets collected from the Internet, it is inappropriate to evaluate LLMs’ data-fitting capabilities on publicly available datasets which may have already been exposed to the LLMs during training. Therefore, we create synthetic data from scratch for all experiments in this paper. Specifically, we simulate the following linear data generation process for $p$ input variables:
@@
In the baseline setting, $p=10$ and each input variable $X_{i}$ is randomly and independently sampled from either $N(0,1)$ or $\exp(1)$; coefficients $\{\beta_{1},\ldots,\beta_{p}\}$ are sampled from $Unif[0,1]$ and fixed; and $\varepsilon$ is randomly sampled from $N(0,0.1^{2})$. We choose such a simple linear data-generation process for two reasons. First, its parametric form gives us fine-grained control over different aspects of the data-fitting task, such as number of input variables ($p$) and distributional characteristics of these variables. Second, it represents an “easy” data-fitting task. If LLMs exhibit problematic data-fitting capabilities under this setup, it is more likely due to a fundamental limitation of current LLMs rather than difficulty of the data-fitting task. We use this data-generation process to simulate a training dataset of 4,000 instances and a testing dataset of 1,000 instances.
@@
We focus on two specific data-fitting capabilities. We first evaluate the performance level, i.e., how well LLMs fit data and generate predictions compared to tabular supervised learning techniques. This determines whether LLMs have the capability to fit data accurately. More importantly, we also evaluate task-irrelevant prediction sensitivity, i.e., whether LLM-generated predictions are robust with respect to changes in the inputs that are irrelevant to the data-fitting tasks (and are extraneous for tabular supervised learning techniques). This sheds light on whether LLMs should be used for data fitting.
@@
We carry out comprehensive data-fitting experiments on GPT-4o-mini (a representative general-purpose close-weight LLM), [^3] Among many close-weight commercial LLMs, GPT-4o-mini is one of the most cost-effective flagship LLMs that support both completion and fine-tuning. Llama-3-8B-instruct (a representative general-purpose open-weight LLM), as well as TabPFN (currently best-performing tabular foundation model specifically trained for data-fitting tasks, Hollmann et al.,, 2025). We perform LLM-based data fitting via both in-context learning (ICL) and supervised fine-tuning (SFT). For GPT-4o-mini, we use OpenAI’s batch processing API for ICL and fine-tuning API for SFT. For Llama-3-8B-instruct, we download the model from Hugging Face [^4] Source: https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct. and carry out ICL using resources provided by Minnesota Supercomputing Institute. For general-purpose LLMs (GPT-4o-mini and Llama-3-8B-instruct), we follow standard practices in LLM prompting to represent data instances in a natural language format (e.g., ”A data point has X0 [VAL]…”). We also consider an alternative data format (i.e., JSON) as one task-irrelevant variation. The detailed prompts and procedures for both ICL and SFT are reported in SI Appendix, “LLM Prompts and Prediction Procedures” Section. For the tabular foundation model (TabPFN), we use its official Python implementation to perform data fitting via ICL. Given the objective nature of the task, we use greedy decoding for all LLMs (i.e., selecting the next token to be the one receiving highest predicted probability among the candidate tokens). This is done by setting LLM generation temperature to 0 and disabling random token sampling whenever possible. We also fix the random seed during SFT of LLMs as well as the training of all tabular supervised learning models to ensure reproducibility and fair comparisons across different models. All models are evaluated on the same 1,000 testing instances.
Due to the non-trivial monetary and computational costs of LLM usage, we first perform data fitting and evaluations with GPT-4o-mini and Llama-3-8B-instruct on a single (fixed) set of training and testing data (and report the corresponding results). We then repeat data fitting with these LLMs for a second time, in order to test whether the observed prediction sensitivity is caused by task-irrelevant variations or by some inherent randomness in LLMs’ token generation. Furthermore, for analyses involving TabPFN, since the tabular foundation model can be run locally and efficiently, we repeat the data-fitting experiments 100 times to obtain statistical evidence of prediction sensitivity.
Replication materials can be accessed at https://github.com/heziiiiiiiii/LLMs-for-data-fitting.
@@ 3.1 Performance Level
Starting with GPT-4o-mini, we evaluate the LLMs’ predictive performance level as compared to six widely-used tabular supervised learning techniques, including linear regression (the oracle model in this case), LASSO regression, $k$-NN, Support Vector Regression (SVR), Random Forest, and Multi-Layer Perceptron (MLP). Each supervised technique is trained on the same training data of 4,000 instances, with hyper-parameters tuned via 5-fold cross-validation. For reference, we also include a naive benchmark, namely using the mean of outcome variable on the training data as (constant) prediction. For GPT-4o-mini under SFT, we use all 4,000 training instances for fine-tuning. Under ICL, following the common practice in few-shot prompting, we include either 10 or 20 instances sampled from training data into the task prompt (denoted as ICL10 and ICL20). We also explore an ICL prompt with 500 instances (denoted as ICL500), representing a large-scale setting that is still within the context window of GPT-4o-mini. We evaluate predictive performance on the testing data using the standard Mean Absolute Error (MAE) metric. The results are shown in Table 1.
@@
Not surprisingly, linear regression and other tabular supervised learning techniques that can sufficiently approximate a linear relationship (LASSO, SVR, and MLP) all perform quite well on this task, whereas Random Forest and $k$-NN perform comparatively worse. More importantly, LLM-based data fitting achieves respectable performance. With SFT, LLM outperforms Random Forest and $k$-NN and comes somewhat close to the performance of the other techniques. With ICL, even with only 10 or 20 examples (substantially smaller “training” size than the other models), it still meaningfully outperforms the naive benchmark, indicating that LLM can conduct few-shot learning to some degree in a prediction task. Further increasing the number of examples in ICL to 500 does not necessarily improve predictive performance. [^5] We further consider a few variations to the above baseline setting, including different numbers of variables, variable scales, and the presence of outliers in data. Overall, we find that LLMs can achieve reasonable performance levels across these variations, especially with SFT.